# AdaCQR: Enhancing Query Reformulation for Conversational Search via Sparse and Dense Retrieval Alignment

### v1

@@ Abstract
Conversational Query Reformulation (CQR) has significantly advanced in addressing the challenges of conversational search, particularly those stemming from the latent user intent and the need for historical context. Recent works aimed to boost the performance of CRQ through alignment. However, they are designed for one specific retrieval system, which potentially results in poor generalization. To overcome this limitation, we present a novel framework AdaCQR. By aligning reformulation models with both term-based and semantic-based retrieval systems, AdaCQR enhances the generalizability of information-seeking queries across diverse retrieval environments through a two-stage training strategy. We also developed two effective approaches for acquiring superior labels and diverse input candidates, boosting the efficiency and robustness of the framework. Experimental evaluations on the TopiOCQA and QReCC datasets demonstrate that AdaCQR significantly outperforms existing methods, offering both quantitative and qualitative improvements in conversational query reformulation.111  Code and Datasets: https://github.com/init0xyz/AdaCQR
@@
Yilong Lai   Jialong Wuthanks:   Equal Contribution.   Congzhi Zhang   Haowen Sun   Deyu Zhouthanks:   Corresponding Author.  School of Computer Science and Engineering, Southeast University, Nanjing, China Key Laboratory of New Generation Artificial Intelligence Technology and Its Interdisciplinary Applications (Southeast University), Ministry of Education, China {yilong.lai, jialongwu, zhangcongzhi, haowensun, d.zhou}@seu.edu.cn
@@ 1 Introduction
Conversational search extends traditional information retrieval paradigms by addressing complex information-seeking requirements through multi-turn interactions  (Radlinski and Craswell, 2017; Qu et al., 2020; Gao et al., 2023). A fundamental challenge in conversational search is to discover the latent user intent within the current query and historical context, which complicates the application of off-the-shelf retrievers due to issues such as omissions, ambiguity, and coreference (Anantha et al., 2021; Adlakha et al., 2022).
Existing methods to address this challenge can be broadly categorized into two types: dense retriever-based and query reformulation-based. For dense retrievers-based approaches (Qu et al., 2020; Lin et al., 2021b; Kim and Kim, 2022; Mo et al., 2024; Chen et al., 2024), long dialogue contexts can be effectively grasped while incurring retraining costs and lacking the adaptability to sparse retrieval systems like BM25 (Robertson et al., 2009). Query reformulation-based approaches leverage a language model to decontextualize the query of user into a stand-alone query, a process known as conversational query reformulation (CQR), as shown in Figure 1. Previous studies have demonstrated the effectiveness of CQR Wu et al. (2022); Mo et al. (2023a); Ye et al. (2023).
@@
Due to the limitation that the training objectives do not align with task targets, i.e., minimizing cross-entropy loss for teacher forcing generation during training while expecting to maximize retrieval metric during inference, subsequent works have aimed to enhance the performance of CQR through alignment. In detail, Jang et al. (2023) utilize Minimum Bayes Risk (MBR) Smith and Eisner (2006) based on semantic similarity between the query and gold passage to achieve alignment. Yoon et al. (2024) create binarized comparisons based on retriever feedback and optimize the reformulation model via Direct Preference Optimization (DPO) Rafailov et al. (2023). They also tackle the reliance on sub-optimal and costly human-annotated reformulation labels by using Large Language Models (LLMs) to generate labels via iterative prompting or multi-perspective prompting.
@@
However, previous alignment methods are designed for one specific retrieval system, which may fail to be generalized to multiple retrieval systems simultaneously. For an information-seeking query to generalize well across both sparse and dense retrieval systems, it must have: (1) precise term overlap (e.g., the presence of key entities in the query) and (2) high semantic similarity between the document and the query Luan et al. (2021). Focusing on only one of these aspects would lead to performance degradation. In addition, previous works for achieving alignment exhibit stability issues when using reinforcement learning  Jang et al. (2023) and require an explicit reference model  Jang et al. (2023); Yoon et al. (2024).
@@
Therefore, in this paper, we introduce AdaCQR, a novel framework that aligns the training objective with the task target. In specific, AdaCQR aligns the reformulation model and the retrievers from both term-based and semantic-based perspectives to achieve strong generalization abilities in sparse and dense retrieval. Furthermore, to address the issues of high complexity and instability inherent in MBR  Jang et al. (2023), we employ a two-stage training strategy to achieve alignment Liu et al. (2022), where the reformulation model serves both as a generation model using cross-entropy loss for teacher forcing generation and a reference-free evaluation model using contrastive loss.
The framework works as follows: 1) An advanced generation model is initialized with superior reformulation labels. Specifically, a few-shot LLM prompting method is employed inspired by the principles of contrastive learning Paranjape et al. (2021); He et al. (2022) to get superior labels; 2) Unlike previous methods that rely on human or LLMs annotations, Diverse Beam Search Vijayakumar et al. (2016) is used to generate multiple candidates simultaneously. Among the generated candidates, one oracle candidate exhibits exceptional performance, while the remaining candidates are relatively ranked based on a fusion metric; 3) We employ a two-stage training, where the reformulation model can be aligned using the contrastive loss from both term-based and semantic-based perspectives.
@@
AdaCQR achieves excellent performance on two widely used conversation search datasets, TopiOCQA  Adlakha et al. (2022) and QReCC  Anantha et al. (2021). Notably, AdaCQR achieves the performance comparable to those approaches fine-tuned on the LLaMA-7B backbone, despite of utilizing only the T5-base. Experimental results demonstrate the quantitative and qualitative improvements of our proposed framework.
@@
  • We propose AdaCQR to align reformulation models with term-based and semantic-based retrievers, simultaneously.
  • Two effective approaches are developed: one to acquire superior labels for generation and another to gather diverse input candidates for reference-free evaluation.
  • Extensive experiments on two benchmark datasets conclusively demonstrate our proposed AdaCQR significantly outperforms existing methods, establishing its superiority in performance.
@@ 2.1 Conversational Search
Conversational search improves traditional information retrieval by using iterative, multi-turn interactions to address the complex information needs of user  Gao et al. (2023). A key challenge is understanding the the implicit intent of user, requiring attention to both the current query and its historical context. Two main approaches to this problem are conversational dense retrieval (CDR) and conversational query reformulation (CQR).
CDR  Qu et al. (2020); Yu et al. (2021); Lin et al. (2021b) aims to improve the representation of the current query along with its historical context by training dense retrievers. Recent advancements in CDR have focused on mitigating the influence of irrelevant historical contexts  Kim and Kim (2022); Mo et al. (2023b, 2024); Chen et al. (2024) and enhancing interpretability  Mao et al. (2023b); Cheng et al. (2024). However, this approach incurs additional training costs and lacks the adaptability to sparse retrieval systems like BM25  Robertson et al. (2009).
Conversely, CQR  Elgohary et al. (2019) concentrates on decontextualizing the query of user into a stand-alone query suitable for use with off-the-shelf retrievers. Numerous prior studies have demonstrated the effectiveness of CQR by utilizing human annotations in supervised methods  Lin et al. (2020a); Yu et al. (2020); Vakulenko et al. (2021) and integrating query expansion models Mo et al. (2023a). However, human-annotated labels are costly and reported to be sub-optimal Lin et al. (2021b); Wu et al. (2022). In the era of LLMs, several studies have utilized LLMs to generate query reformulations directly Ye et al. (2023); Mao et al. (2023a) and obtain reformulation labels for distillation Jang et al. (2023); Yoon et al. (2024).
This paper focuses on conversational query reformulation, proposing a novel framework AdaCQR to align with term-based and semantic-based retrieval systems. To overcome the limitations of human annotation, we also developed two effective methods for obtaining superior labels and diverse input candidates.
@@ 2.2 Aligning LMs using Feedback
Aligning language models with feedback involves adjusting their behavior and outputs based on evaluation feedback Wang et al. (2023), employing various reward learning methodologies to provide accurate supervised signals Schulman et al. (2017); Rafailov et al. (2023).
Recent studies have enhanced conversational query reformulation by aligning language models with retriever feedback Jang et al. (2023); Yoon et al. (2024). Jang et al. (2023) achieve the alignment through minimizing Bayes Risk based on semantic similarity between the query and the gold passage. Yoon et al. (2024) leverages LLMs to generate numerous reformulations via multi-perspective prompting, creating binarized comparisons based on retriever feedback and optimizing the reformulation model using DPO  Rafailov et al. (2023). However, previous methods struggle with the high cost of generating reformulations with LLMs Yoon et al. (2024), or the instability of MBR Jang et al. (2023); Finkelstein and Freitag (2023).
In contrast, our framework utilizes a contrastive loss  Liu et al. (2022) to achieve alignment with retrievers. To the best of our knowledge, we are the first to employ the language model as a reference-free evaluation model to align retrievers, thereby enhancing stability and reducing complexity.
@@ 3.1 Task Formulation
The conversational search task discussed in this paper involves finding the passage most relevant to the intent of user from a large collection of passages $\mathcal{C}$, given the current query of the user and historical context. To achieve this goal, the CQR task is proposed to utilize a language model $G_{\theta}$ to condense the current query $q_{k}$ and historical context $\mathrm{H}_{k-1}=\{q_{i},r_{i}\}_{i=1}^{k-1}$ into a stand-alone query $\hat{Q}_{k}$, where $q_{i}$ and $r_{i}$ denote the query and system answer of the $i$-th turn conversation, with $k$ indicating the current turn. This decontextualized query $\hat{Q}_{k}$ is subsequently input into an off-the-shelf retrieval system $\mathcal{R}$, which returns a ranked list of the top-$k$ relevant passages.
For the sake of convenience, we define the CQR task as the problem $\mathcal{P}=\{q,\mathrm{H}\}$, where $q$ represents the current query of user and $\mathrm{H}$ denotes the historical context. The task is to generate a reformulated query $\hat{Q}$, as discussed in the following sections.
@@ 3.2 Overall framework
Our framework begins to leverage LLM to generate superior reformulation labels $Q^{\star}$ via few-shot learning (§3.4), where we select representative examples and implicitly guide LLM to generate labels that meet the needs of retrievers. Subsequently, we employ a two-stage training strategy using these labels to align the reformulation model with the retrievers. In the first stage, we train the reformulation model with a cross-entropy loss $\mathcal{L}_{g}$ to acquire the basic ability to generate reformulation queries using the superior $Q^{\star}$. (§3.5.1). Afterwards, we use this model to create a diverse set $\mathcal{S}$, including candidate queries $C_{(1)},\cdots,C_{(n)}$3.5.2). These candidates are then evaluated through sparse and dense retrieval, assessing their performance from both term-based and semantic-based perspectives. We utilize a proposed fusion metric (§3.3) to synthesize these evaluations and obtain the relative order of the candidates. In the second stage, leveraging the relative order of the candidates, we apply a contrastive loss $\mathcal{L}_{c}$3.5.3) to achieve alignment between the reformulation model and the retrievers, where the reformulation model is treated as an evaluation model. The overall framework is depicted in Figure 2.
@@ 3.3 Fusion Metric for Sparse and Dense Retrieval
To measure the generalization ability of the reformulation queries, we input them into sparse and dense retrieval systems and assess their performance based on the ranking of the corresponding gold passages, as illustrated in the central part of Figure 2.
In sparse retrieval, the inverted index is constructed using the sparse vectors of the transformed documents. The query is then tokenized into terms and matches passage based on term overlap. In contrast, dense retrieval involves creating a vectorized index using the dense vectors of the transformed documents. Subsequently, the query is converted into an embedding by the encoder, and the passage is searched based on semantic similarity.
Leveraging the performance of the reformulation query in both retrieval systems, we propose a fusion metric to evaluate the performance of reformulation query $\hat{Q}$ more comprehensively, similar to reciprocal rank fusion  Cormack et al. (2009):
$\mathrm{M}(\hat{Q},d)=\frac{r_{s}(\hat{Q},d)+r_{d}(\hat{Q},d)}{r_{s}(\hat{Q},d)\times r_{d}(\hat{Q},d)}$ (1)
where $\hat{Q}$ is a reformulation query, $d$ is the gold passage. $r_{s}(q,d)$ and $r_{d}(q,d)$ represent the rank of the gold passage $d$ within the sparse and dense retrieval results for query $q$, respectively. The ranking $r_{s}$ and $r_{d}$ starts from $1$, indicating the highest-ranked passage. Based on Eq (1), a larger $\mathrm{M}(\hat{Q},d)$ indicates better generalization performace for reformulation query $\hat{Q}$ on sparse and dense retrieval systems.
@@ 3.4 Superior Reformulation Annotation
To mitigate the dependency on costly and sub-optimal human-annotated reformulation labels, we utilize LLMs to generate superior reformulation labels, offering a more robust foundation for our framework.
Our intuitive approach to obtaining superior reformulation labels is to convey the LLMs of the characteristics of an effective query reformulation for retrieval. However, defining a good reformulation query or providing explicit instructions for generating one that meets the needs of retrievers poses significant challenges. Leveraging excellent in-context learning capabilities of LLMs Brown et al. (2020); Dong et al. (2022), we propose a prompting strategy that implicitly selects representative demonstrations and guides LLMs to generate reformulation labels aligned with the requirements of retrieval systems.
Our method begins with a vanilla generative model $G_{\pi}$ with basic query reformulation ability. We employ $G_{\pi}$ to generate a reformulation candidates set $\mathcal{S}_{\pi}=\{C_{(1)},C_{(2)},\cdots,C_{(n)}\}$ for each the reformulation problem $\mathcal{P}=\{q,\mathrm{H}\}$, on the validation set. To select representative demonstrations, we use a score $\hat{R}$ to describe the difficulty of the query reformulation problem $\mathcal{P}$:
$\hat{R}=\frac{1}{n}\sum_{i=1}^{n}[\mathrm{M}(C_{(i)},d)-\frac{1}{n}\sum_{j=1}^{n}\mathrm{M}(C_{(j)},d)]^{2}$ (2)
where $C_{(i)}$ is a reformulation candidate generated by $G_{\pi}$, and $d$ is the corresponding gold passage for reformulation problem $\mathcal{P}$. The metric $\mathrm{M}(\cdot)$ is defined according to Eq. (1).
Subsequently, we selected the top-$m$ reformulation problems exhibiting the highest $\hat{R}$ scores from the validation set. For each selected reformulation problem $\mathcal{P}=\{q,\mathrm{H}\}$, we identified the best and worst reformulation candidates from set $\mathcal{S}_{\pi}$ based on the metric in Eq. (1). These candidates denoted as $C_{\text{best}}$ and $C_{\text{worst}}$, serve to implicitly guide the LLMs in generating labels aligned with the needs of the retrieval system, inspired by contrastive learning  Paranjape et al. (2021); He et al. (2022). We then concatenate the demonstration $P=(q,\mathrm{H},C_{\text{best}},C_{\text{worst}})$ and task instruction $\mathcal{I}$ to form the final prompt $\mathcal{D}=\mathcal{I}\ ||P_{1}||\cdots||P_{m}$, where ||||| | donates concatenation. Finally, we employed the LLM to obtain the superior reformulation labels $Q^{\star}$ through in-context learning. The details for the annotation are presented in Appendix D.

### v3

@@ Abstract
Conversational Query Reformulation (CQR) has significantly advanced in addressing the challenges of conversational search, particularly those stemming from the latent user intent and the need for historical context. Recent works aimed to boost the performance of CQR through alignment. However, they are designed for one specific retrieval system, which potentially results in sub-optimal generalization. To overcome this limitation, we present a novel framework AdaCQR. By aligning reformulation models with both term-based and semantic-based retrieval systems, AdaCQR enhances the generalizability of information-seeking queries among diverse retrieval environments through a two-stage training strategy. Moreover, two effective approaches are proposed to obtain superior labels and diverse input candidates, boosting the efficiency and robustness of the framework. Experimental results on the TopiOCQA, QReCC and TREC CAsT datasets demonstrate that AdaCQR outperforms the existing methods in a more efficient framework, offering both quantitative and qualitative improvements in conversational query reformulation.111[Uncaptioned image] : https://github.com/init0xyz/AdaCQR
@@
Yilong Laithanks:   Equal Contribution.,   Jialong Wu,   Congzhi Zhang,   Haowen Sun,   Deyu Zhouthanks:   Corresponding Author.  School of Computer Science and Engineering, Key Laboratory of Computer Network and Information Integration, Ministry of Education, Southeast University, China {yilong.lai, jialongwu, zhangcongzhi, haowensun, d.zhou}@seu.edu.cn
@@ 1 Introduction
Conversational search extends the traditional information retrieval paradigms by addressing complex information-seeking requirements through multi-turn interactions  (Radlinski and Craswell, 2017; Qu et al., 2020; Gao et al., 2023). A fundamental challenge in conversational search is to discover the latent user intent within the current query and historical context, which complicates the application of off-the-shelf retrievers due to issues such as omissions, ambiguity, and coreference (Anantha et al., 2021; Adlakha et al., 2022).
Existing methods to address this challenge can be broadly categorized into two types: dense retriever-based and query reformulation-based. For dense retriever-based approaches (Qu et al., 2020; Lin et al., 2021b; Kim and Kim, 2022; Mo et al., 2024d), long dialogue contexts can be effectively grasped by the dense retriever while incurring retraining costs and lacking the adaptability to sparse retrieval systems like BM25 (Robertson et al., 2009). Query reformulation-based approaches leverage a language model to decontextualize every user’s query into a stand-alone query, a process known as conversational query reformulation (CQR), as shown in Figure 1. Previous studies have demonstrated the effectiveness of CQR Wu et al. (2022); Mo et al. (2023a); Ye et al. (2023).
@@
As the training objectives are not aligned with task targets, i.e., minimizing the cross-entropy loss for teacher forcing generation during training while expecting to maximize retrieval metric during inference, subsequent approaches have aimed to enhance the performance of CQR through alignment. For example, Jang et al. (2023) utilize Minimum Bayes Risk (MBR) Smith and Eisner (2006) based on semantic similarity between the query and gold passage to achieve alignment. Yoon et al. (2024) create binarized comparisons based on retriever feedback and optimize the reformulation model via direct preference optimization. They also tackle the reliance on sub-optimal and costly human-annotated reformulation labels by using Large Language Models (LLMs) to generate labels via iterative prompting or multi-perspective prompting.
@@
However, the methods mentioned above are designed for one specific retrieval system. For an information-seeking query to be generalized well across both sparse and dense retrieval systems, it must have: (1) precise term overlap (e.g., the presence of key entities in the query) and (2) high semantic similarity between the document and the query Luan et al. (2021). Both characteristics of the information-seeking query play a crucial role in query reformulation; jointly leveraging them is beneficial. In addition, previous works leverage reinforcement learning to achieve alignment, but they exhibited stability issues and relied on an explicit reference model  Wu et al. (2022); Jang et al. (2023).
@@
Therefore, in this paper, we introduce AdaCQR, a novel framework that effectively aligns the training objective with the task target. In specific, AdaCQR aligns the reformulation model and the retrievers from both term and semantic perspectives to achieve strong generalization abilities in sparse and dense retrieval. Furthermore, to address the issues of high complexity and instability inherent in MBR, we employ a two-stage training strategy to achieve alignment, where the reformulation model serves both as a generation model using cross-entropy loss for teacher forcing generation and a reference-free evaluation model using contrastive loss.
The framework works as follows: 1) A fusion metric is introduced to evaluate the generalization performance of the reformulated query across various retrieval systems. 2) By leveraging uncertainty estimation derived from the fusion metric, combined with insights from contrastive learning, we implicitly guide the large language model (LLM) to generate superior reformulation labels, which are then utilized to initialize the reformulation model during the first training stage. 3) In the second training stage, Diverse Beam Search Vijayakumar et al. (2016) is employed to effectively gather multiple candidate reformulation queries. The reformulation model is then aligned using a contrastive loss from both term and semantic perspectives, incorporating the candidates and their relative orders based on the fusion metric.
@@
AdaCQR achieves excellent performance on widely used conversation search datasets, including TopiOCQA  Adlakha et al. (2022), QReCC  Anantha et al. (2021) and TREC CAsT Dalton et al. (2020, 2021, 2022). Notably, to maintain the smoothness of the overall system, we selected a lightweight model as our backbone, which achieved performance comparable to those approaches fine-tuned on the LLaMA-7B backbone or aggregated multiple candidate queries from proprietary LLM. Experimental results demonstrate the quantitative and qualitative improvements of our proposed framework.
@@
  • We propose AdaCQR to align reformulation models from both term and semantic perspectives.
  • By leveraging the proposed fusion metric, we can effectively acquire superior labels for generation and collect diverse ordered candidate queries for reference-free evaluation.
  • Extensive experiments on several benchmark datasets conclusively demonstrate our proposed AdaCQR significantly outperforms existing methods, establishing its superiority in performance.
@@ 2.1 Conversational Search
Conversational search improves traditional information retrieval by using iterative, multi-turn interactions to address the complex information needs of users  Gao et al. (2023); Mo et al. (2024b). A key challenge is understanding the implicit intent of the user, requiring attention to both the current query and its historical context. Two main approaches to this problem are conversational dense retrieval (CDR) and conversational query reformulation (CQR).
CDR  Qu et al. (2020); Yu et al. (2021); Lin et al. (2021b) aims to improve the representation of the current query along with its historical context by training dense retrievers. Recent advancements in CDR have focused on mitigating the influence of irrelevant historical contexts  Kim and Kim (2022); Mo et al. (2023b, 2024c, 2024d) and enhancing interpretability  Mao et al. (2023c); Cheng et al. (2024). However, this approach incurs additional training costs and lacks the adaptability to sparse retrieval systems like BM25  Robertson et al. (2009).
Conversely, CQR  Elgohary et al. (2019) concentrates on decontextualizing the query of user into a stand-alone query suitable for use with off-the-shelf retrievers. Numerous prior studies have demonstrated the effectiveness of CQR by utilizing human annotations in supervised methods  Lin et al. (2020); Yu et al. (2020); Vakulenko et al. (2021) and integrating query expansion models Mo et al. (2023a, 2024a). However, human-annotated labels are costly and reported to be sub-optimal Lin et al. (2021b); Wu et al. (2022). In the era of LLMs, several studies have utilized LLMs to generate query reformulations directly Ye et al. (2023); Mao et al. (2023b) and obtain reformulation labels for distillation Jang et al. (2023); Yoon et al. (2024). This paper focuses on conversational query reformulation, proposing a novel framework AdaCQR to align with term-based and semantic-based retrieval systems. To overcome the limitations of human annotation, we also developed two effective methods for obtaining superior labels and diverse input candidates.
@@ 2.2 Aligning LMs using Feedback
Aligning language models with feedback involves adjusting their behaviour and outputs based on evaluation feedback Wang et al. (2023), employing various reward learning methodologies to provide accurate supervised signals Schulman et al. (2017); Rafailov et al. (2023).
Recent studies have enhanced conversational query reformulation by aligning language models with retriever feedback Jang et al. (2023); Yoon et al. (2024). Jang et al. (2023) achieve the alignment through minimizing Bayes Risk based on semantic similarity between the query and the gold passage. Yoon et al. (2024) leverages LLMs to generate numerous reformulations via multi-perspective prompting, creating binarized comparisons based on retriever feedback and optimizing the reformulation model using DPO  Rafailov et al. (2023). However, previous methods struggle with the high cost of generating reformulations with LLMs Yoon et al. (2024), or the instability of MBR Jang et al. (2023); Finkelstein and Freitag (2023). In contrast, our framework utilizes a contrastive loss  Liu et al. (2022) to achieve alignment with retrievers. To the best of our knowledge, we are the first to employ the language model as a reference-free evaluation model to align retrievers, thereby enhancing stability and reducing complexity.
@@ 3.1 Task Formulation
The conversational search task discussed in this paper involves finding the passage most relevant to the intent of the user from a large collection of passages, given the current query of the user and historical context. To achieve this goal, the CQR task is to utilize a language model $G_{\theta}$ to condense the current query $q_{k}$ and historical context $\mathrm{H}_{k-1}=\{q_{i},r_{i}\}_{i=1}^{k-1}$ into a stand-alone query $\hat{Q}_{k}$, where $q_{i}$ and $r_{i}$ denote the query and system answer of the $i$-th turn conversation, with $k$ indicating the current turn. This decontextualized query $\hat{Q}_{k}$ is subsequently input into an off-the-shelf retrieval system $\mathcal{R}$, which returns a ranked list of the top-$p$ relevant passages.
For sake of convenience, we formalize the CQR task as a session $\mathcal{P}=\{q,\mathrm{H}\}$, where $q$ represents the user’s current query and $\mathrm{H}$ denotes the historical context. The objective is to generate a reformulated query $\hat{Q}$, as discussed in the following sections.
@@ 3.2 Overall Framework
Figure 2 depicts the overall framework of AdaCQR. The framework begins to introduce a fusion metric to evaluate the generalization performance of the reformulation queries across sparse and dense retrieval systems(§3.3). Leveraging the uncertainty estimation based on the fusion metric and the insight from contrastive learning, we select representative examples and implicitly guide LLM to generate superior labels $Q^{\star}$ to meet the needs of retrievers(§3.4). Subsequently, we employ a two-stage training strategy to align the reformulation model with the retrievers. In the first stage, we train the reformulation model with a cross-entropy loss $\mathcal{L}_{g}$ using the superior $Q^{\star}$ to acquire the basic ability to generate reformulation queries(§3.5.1). Afterwards, the model is used to create a diverse set $\mathcal{S}$ comprising candidates queries $C_{(1)},\cdots,C_{(n)}$3.5.2), which are then evaluated across sparse and dense retrieval, considering from term and semantic perspectives, and ranked based on a proposed fusion metric that synthesizes these evaluations. In the second stage, leveraging the relative order of the candidates, we apply a contrastive loss $\mathcal{L}_{c}$3.5.3) to achieve alignment between the reformulation model and the retrievers, where the reformulation model acts as an evaluation model.
@@ 3.3 Fusion Metric for Sparse and Dense Retrieval
In sparse retrieval, the query is tokenized into terms and matches passage based on term overlap. In contrast, dense retrieval converts the query into an embedding by the encoder and searches passages based on semantic similarity. To measure the generalization ability of the reformulation queries, we input them into sparse and dense retrieval systems and assess their performance based on the ranking of the corresponding gold passages, as illustrated in the central part of Figure 2.
Inspired by reciprocal rank fusion  Cormack et al. (2009), we propose a fusion metric to merge the score of sparse retrieval based on term overlap and dense retrieval based on semantic similarity into an optimized score:
$\mathrm{M}(\hat{Q},d)=\frac{1}{r_{s}(\hat{Q},d)}+\frac{1}{r_{d}(\hat{Q},d)}$ (1)
where $\hat{Q}$ is a reformulation query, $d$ is the gold passage. $r_{s}(q,d)$ and $r_{d}(q,d)$ represent the rank of the gold passage $d$ within the sparse and dense retrieval results for query $q$, respectively. The ranking $r_{s}$ and $r_{d}$ starts from $1$, indicating the highest-ranked passage.
Leveraging Eq (1), the generalization performance of the reformulation query $\hat{Q}$ could be evaluated, where a larger $\mathrm{M}(\hat{Q},d)$ indicates better generalization performance for reformulation query $\hat{Q}$ on sparse and dense retrieval systems.
@@ 3.4 Superior Reformulation Annotation
We leverage LLMs to generate high-quality reformulation labels by conveying the characteristics of effective query reformulation for retrieval. Given the challenge of explicitly defining optimal reformulation, we propose a prompting strategy that selects representative examples and implicitly guides LLMs to generate reformulation labels based on contrastive learning.
It begins with a vanilla model $G_{\pi}$ with basic query reformulation ability. We employ $G_{\pi}$ to generate a reformulation candidates set $\mathcal{S}_{\pi}$ with diverse beam search for each reformulation session. Motivated by previous works showing that the annotated most uncertain examples can significantly enhance the effectiveness of in-context learning Diao et al. (2024); Yue et al. (2024), the variance $\operatorname{Var}(\mathcal{S}_{\pi})$, is utilized as a measure for estimating uncertainty in reformulation tasks, where scores for ${S}_{\pi}$ are computed using the fusion metric described in Eq (1). A higher variance suggests greater instability in the performance of the retriever, indicating a more challenging reformulation problem. We then selected the top-$m$ representative reformulation problems exhibiting the highest variances on the validation set.
For the representative example annotation, inspired by contrastive learning Paranjape et al. (2021); He et al. (2022), we identified the best and worst reformulation candidates from set $\mathcal{S}_{\pi}$ based on Eq (1), denoted as $C_{\text{best}}$ and $C_{\text{worst}}$ to implicitly guide the LLMs in generating labels aligned with the needs of the retrieval system. We then concatenate the $m$ representative demonstrations, each consists of $(q,\mathrm{H},C_{\text{best}},C_{\text{worst}})$, along with task instruction $\mathcal{I}$. Finally, we employed the LLM to obtain the superior reformulation labels $Q^{\star}$ through in-context learning Brown et al. (2020); Dong et al. (2022); Xiang et al. (2024). The details of the annotation are presented in Appendix D.